in order to decrease radioactive nuclei to one million of its initial number,number of half lives required is

Grade:12th pass

1 Answers

Vikas TU
6 years ago
Dear Student,
For the firrst order of chemcial kinetics,
t=2.303log(a/a-x)/l
=>t=2.303log(a/a-106)/l
no of half lives=t/t1/2=(2.303log(a/a-106)/l)/(0.693/l) (since t1/2=0.693/l)
no of half lives can be found by substituting a which is the initial number that is not given in the question.
Cheers!!
Regards,
Vikas (B. Tech. 4th year
Thapar University)